OpenStudy (anonymous):

simplify: sqrt 20/ sqrt 8

OpenStudy (anonymous):

sqrt(5/2)

OpenStudy (precal):

\[(\sqrt{20})/(\sqrt{8})\] \[(\sqrt{2*2*5})/(\sqrt{2*2*2})\] \[(2\sqrt{5})/(2\sqrt{2})\] \[(\sqrt{5})/(\sqrt{2})\]